describe the shape and position of the graph of y = (x - h)^2 + g.
Answer by AnlytcPhil(1806)

>>...describe the shape and position of the graph of
y = (x - h)^2 + g.
Usually books use the letter "k" where you have "g". Also they
usually have the coefficient "a" in front of the first term
y = a(x - h)² + k
represents the graph of a parabola, a U-shaped graph that has
vertex at (h,k) and passes through the two points (h±1, k+a).
If "a" is a positive number the parabola opens upward with the
vertex at the bottom. If "a" is negative the parabola opens
downward with the vertex at the top.
Yours is the special case where a=1 and k=g
y = (x - h)² + g
which represents the graph of a parabola, a U-shaped graph that
opens upward, has vertex at (h,g) and passes through the two
points (h±1, g+1).
Edwin